OpenStudy (christos):

f(x) = cuberoot(-8x-6) find F^-1(x)

OpenStudy (anonymous):

that means to find the inverse of \(f(x)\)\(\rightarrow~f^{-1}(x)\) that being said, if: \[f(x)=\sqrt[3]{-8x-6}\]replace \(f(x)\) with \(x\), and \(x\) with \(y\), then solve for \(y\):\[x=\sqrt[3]{-8y-6}\]cube both sides to rid of the cuberoot:\[x^3=-8y-6\]add \(6\) to both sides:\[x^3+6=-8y\]divide both sides by \(-8\):\[{{x^3+6}\over{-8}}=y\]now that you've solved for \(y\), replace it with \(f^{-1}(x)\):\[f^{-1}(x)={{x^3+6}\over{-8}}\]and that is how you find the inverse of a function :)

OpenStudy (anonymous):

hope that's helpful! @Christos

OpenStudy (christos):

it really is!!!

OpenStudy (anonymous):

NICE.